# MoVieS: Motion-Aware 4D Dynamic View Synthesis in One Second

### v1

@@ Abstract
We present MoVieS, a novel feed-forward model that synthesizes 4D dynamic novel views from monocular videos in one second. MoVieS represents dynamic 3D scenes using pixel-aligned grids of Gaussian primitives, explicitly supervising their time-varying motion. This allows, for the first time, the unified modeling of appearance, geometry and motion, and enables view synthesis, reconstruction and 3D point tracking within a single learning-based framework. By bridging novel view synthesis with dynamic geometry reconstruction, MoVieS enables large-scale training on diverse datasets with minimal dependence on task-specific supervision. As a result, it also naturally supports a wide range of zero-shot applications, such as scene flow estimation and moving object segmentation. Extensive experiments validate the effectiveness and efficiency of MoVieS across multiple tasks, achieving competitive performance while offering several orders of magnitude speedups.
[…]
@@ 1 Introduction
Humans and animals perceive a continuous stream of observations from a dynamic three-dimensional world, and effortlessly interpret its underlying 3D geometry and motion. Replicating this capability is essential for any embodied agent that must understand and act in the physical world.
@@
Recent advances have made great strides in individual 3D tasks, such as monocular depth estimation (ranftl2020towards, ; ke2024repurposing, ; yang2024depth, ; bochkovskii2024depth, ; chen2025video, ), 3D scene reconstruction (schonberger2016structure, ; teed2021droid, ; wang2024dust3r, ; zhang2024monst3r, ; wang2025vggt, ), novel view synthesis (mildenhall2020nerf, ; fridovich2023k, ; kerbl20233d, ; wu20244d, ; luiten2024dynamic, ) and point tracking (harley2022particle, ; doersch2023tapir, ; wang2023tracking, ; karaev2024cotracker, ; xiao2024spatialtracker, ). However, besides treating each task in isolation, most existing view synthesis and reconstruction studies focus on static scenes and require costly per-scene optimization without learning prior knowledge. Real-world environments are inherently dynamic and diverse, and all the aforementioned 3D scene understanding tasks could share common underlying principles.
@@
Motivated by this, we introduce MoVieS, a Motion-aware dynamic View Synthesis model for feed-forward 4D reconstruction of videos in the wild, that jointly models scenes’ appearance, geometry and motion. MoVieS represents 3D dynamic scenes using renderable and deformable 3D particles, termed dynamic splatter pixels, and utilizes a differentiable 3D Gaussian rendering framework (kerbl20233d, ). Specifically, following recent practices on feed-forward view synthesis (charatan2024pixelsplat, ; szymanowicz2024splatter, ; zhang2024gs, ; xu2024depthsplat, ; lin2025diffsplat, ), each input pixel is mapped to a 3D Gaussian primitive, with its 3D location determined by predicted depth. To model dynamics, MoVieS regresses per-pixel motion displacements toward arbitrary query timestamps, enabling temporal tracking of each splatter pixel. This design facilitates coherent reconstruction of both 3D geometry and appearance across camera viewpoints and temporal frames.
@@
As shown in Figure 1, MoVieS is build upon a large-scale pretrained transformer backbone (oquab2024dinov2, ; wang2025vggt, ), which encodes each video frame independently and aggregates their information via attentions (vaswani2017attention, ). The aggregated features are then processed by specialized prediction heads: (1) a depth head estimates depth for each input frame, (2) a splatter head predicts per-pixel 3D Gaussian (kerbl20233d, ) appearance attributes, such as color and opacity, for novel view rendering, (3) a motion head estimates the time-conditioned movements of Gaussian primitives towards a target timestamp, allowing us to track its temporal evolution.
@@
Benefiting from the unified architecture, MoVieS can be trained on large-scale datasets featuring both static (zhou2018stereo, ; wang2020tartanair, ; li2023matrixcity, ) and dynamic (mehl2023spring, ; cabon2020virtual, ) scenes, as well as point tracking datasets (zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ). At inference time, MoVieS takes a monocular video, whether depicting a static or dynamic scene, and reconstructs per-pixel 3D Gaussian primitives along with their motion attributes at any target timestamp, enabling novel view synthesis, depth estimation, and 3D point tracking in a single model.
@@
Extensive experiments on diverse benchmarks (zhou2018stereo, ; yoon2020novel, ; gao2022dynamic, ; koppula2024tapvid, ) show that MoVieS achieves competitive performance across a variety of 4D perception tasks, while being several orders of magnitude faster than the existing state of the art. Furthermore, empowered by novel view synthesis as a proxy task, MoVieS enables dense motion understanding from sparse tracking supervision. This naturally gives rise to a variety of zero-shot applications, further broadening the potential of our approach.
@@
  • We introduce MoVieS, the first feed-forward framework that jointly models appearance, geometry and motion for 4D scene perception from monocular videos.
@@
Figure 1: Overview. MoVieS consists of a shared image encoder, an attention-based feature backbone (Section 3.2.1), and three heads (Section 3.2.2) that simultaneously predict appearance, geometry and motion. Image shortcut for splatter head and time-varying Gaussian attributes are omitted for brevity.
@@ 2.1 Feed-forward 3D Reconstruction
Most traditional 3D reconstruction methods require dense multi-view supervision and per-scene optimization from scratch at test time (hartley2003multiple, ; davison2007monoslam, ; agarwal2011building, ; mur2015orb, ; schonberger2016structure, ; mildenhall2020nerf, ; kerbl20233d, ). Recent learning-based approaches leverage large-scale data priors to enable feed-forward 3D reconstruction that directly regresses geometric structures such as camera poses and depth (wang2024vggsfm, ; wang2025moge, ; ranftl2020towards, ; piccinelli2024unidepth, ; bochkovskii2024depth, ; yang2024depth, ; hu2024metric3d, ). With the advent of differentiable rendering techniques like 3D Gaussian Splatting (3DGS) (kerbl20233d, ), several works propose to predict pixel-aligned Gaussian attributes for novel view synthesis (charatan2024pixelsplat, ; chen2024mvsplat, ; zhang2024gs, ; ziwen2024long, ; xu2024depthsplat, ). We build on these advances and extend them to dynamic 3DGS representations for novel timestamps.
DUSt3R (wang2024dust3r, ) pioneered direct regression of pixel-aligned pointmaps in a canonical space from image pairs. Subsequent works improved this framework by incorporating feature matching (leroy2024grounding, ), supporting multi-view inputs (wang2025d, ; yang2025fast3r, ; zhang2025flare, ; tang2024mv, ), adapting to streaming video (liu2024slam3r, ; murai2024mast3r, ), or bridging view synthesis with 3DGS (smart2024splatt3r, ; ye2024no, ; zhang2025flare, ; tang2024mv, ). VGGT (wang2025vggt, ) further advances this line by using a stronger image encoder (oquab2024dinov2, ) and multiple task-specific heads to predict all key 3D properties from as few as one to hundreds of views. We adopt VGGT as the backbone of MoVieS for its strong visual representations. However, all these methods are restricted to static scenes and overlook dynamic environments with moving objects—the main focus of our work.
2.2 Dynamic Geometry Reconstruction
Compared to static scene reconstruction, which has been studied for decades, dynamic scene reconstruction remains relatively underexplored. Some works extend the DUSt3R framework to dynamic settings in a plug-and-play manner (chen2025easi3r, ), or leverage foundation models (zhang2024monst3r, ; lu2024align3r, ; yao2025uni4d, ) by incorporating monocular depth (bochkovskii2024depth, ; piccinelli2025unidepthv2, ), optical flow (teed2020raft, ), or 2D point tracking (karaev2024cotracker, ). Dynamic Point Map (sucar2025dynamic, ), a concurrent work, proposes a time-invariant DUSt3R variant by predicting each viewpoint’s pointmap at another timestamp. The most related work, DynaDUSt3R (jin2024stereo4d, ), adds pixel-level motion supervision to model dynamic geometry. However, all these methods are limited to two-frame inputs and produce only point clouds, without high-quality visual reconstructions.
To go beyond two views, CUT3R (wang2025continuous, ) uses a recurrent model to incrementally update scene representations from streaming video for static and dynamic scenes. Other methods (xu2025geometrycrafter, ; jiang2025geo4d, ) treat dynamic reconstruction as conditional video generation, fine-tuning diffusion models to estimate pointmaps probabilistically. However, these require multiple network passes per video.
2.3 Dynamic Novel View Synthesis
Instead of geometric structure, dynamic novel view synthesis emphasizes visual results and view-dependent effects. Early attempts utilize neural radiance field (mildenhall2020nerf, ; fridovich2023k, ; cao2023hexplane, ) to represent dynamic scenes, either from multi-view videos (bansal20204d, ; wang2022fourier, ; li2022neural, ), or monocular ones (pumarola2021d, ; park2021nerfies, ; li2023dynibar, ; zhao2024pgdvs, ). 3DGS (kerbl20233d, ), as a more efficient renderable representation, has been applied for dynamic view synthesis using 4D primitives with a time dimension (yang2024real, ; li2024spacetime, ) or an additional deformable field (wu20244d, ; luiten2024dynamic, ; yang2024deformable, ; lin2024gaussian, ; liu2025modgs, ). Leveraging the low-dimensional nature of scene motion, several works (wang2024shape, ; sun2024splatter, ; stearns2024dynamic, ; lei2025mosca, ; lin2025omniphysgs, ) express motion explicitly and learn Gaussian trajectories across monocular frames. However, these methods are trained from scratch and require iterative optimization for scene adaptation. Off-the-shelf point tracking (doersch2023tapir, ; karaev2024cotracker, ) or optical flow (teed2020raft, ) models are also essential to provide supervision signals.
@@ 2.3 Dynamic Novel View Synthesis
The most related works to ours are BTimer (liang2024feed, ) and NutWorld (shen2025seeing, ), both of which utilize a feed-forward approach to estimate 3DGS attributes from monocular videos. However, BTimer predicts independent Gaussian chunks per timestamp without modeling frame relations and needs an enhancer module for smooth intermediate frames. NutWorld models Gaussian motion but lacks explicit supervision, relying heavily on pretrained depth (chen2025video, ) and flow (xu2023unifying, ), and uses an orthographic camera, which may further lead to projection distortions.
@@ 3 Method
Following previous studies on monocular 4D reconstruction (zhao2024pgdvs, ; wang2024shape, ; liu2025modgs, ; lei2025mosca, ; liang2024feed, ), given a posed video $\mathcal{V}\coloneqq\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i}\}_{i=1}^{N}$ with frames $\mathbf{I}_{i}\in\mathbb{R}^{3\times H\times W}$, camera poses $\mathbf{P}_{i}\in\mathbb{SE}(3)$, intrinsics $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ and timestamps $t_{i}\in[0,1]$, we aim to train a generic model that jointly reconstruct the underlying appearance, geometry and motion of dynamic scenes.
@@ 3.1 Dynamic Splatter Pixel
To model 3D scenes with moving contents, we propose a novel representation, namely dynamic splatter pixel, which decomposes dynamic scenes into a set of static Gaussian primitives and their corresponding deformation fields. Given an input video $\mathcal{V}$, each pixel in the $i$-th frame $\mathbf{I}_{i}$ is associated with a splatter pixel $\mathbf{g}_{i}$ (charatan2024pixelsplat, ; szymanowicz2024splatter, ; zhang2024gs, ) in a shared canonical space of the first frame camera’s coordinate system. Each $\mathbf{g}\coloneqq\{\mathbf{x},\mathbf{a}\}$ is parameterized by its position $\mathbf{x}\in\mathbb{R}^{3}$ in the canonical space and other rendering attributes $\mathbf{a}\in\mathbb{R}^{11}$, including rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$, scale $\mathbf{s}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, and color $\mathbf{c}_{\text{rgb}}\in\mathbb{R}^{3}$ (kerbl20233d, ). Considering splatter pixels were originally designed for static scenes, we decouple motion from geometric structure to adapt them for dynamic scenes. An additional time-dependent deformation field is introduced, in which each splatter pixel $\mathbf{g}$ is associated with $\mathbf{m}(t)\coloneqq\{\Delta\mathbf{x}(t),\Delta\mathbf{a}(t)\}$. $\Delta\mathbf{x}(t)\in\mathbb{R}^{3}$ is the motion vector of a splatter pixel at time $t$ with respect to the canonical 3D space, and $\Delta\mathbf{a}(t)$ is the change of the corresponding attributes of the splatter pixel at time $t$. Therefore, the splatter pixel $\mathbf{g}$ is deformed to time $t$ as:
@@
By combining static splatter pixels and their deformation fields, we establish the correspondence between each Gaussian primitive and its temporal dynamics, thereby enabling dynamic scene modeling and dense motion estimation. In practice, we find that making only the scale and rotation attributes time-dependent, i.e., $\Delta\mathbf{a}(t)\in\mathbb{R}^{7}$, is sufficient for representing dynamic splatter pixels.
Figure 2: Motion Head. Given $t_{q}$ target timesteps, the proposed motion head is conditioned via adaptive layer normalization (AdaLN) and predicts 3D displacements for each input pixel. After rasterization using the $M$ corresponding query-time cameras, output images in shape $M\times 3\times H\times W$ are rendered for supervision. Gaussian attribute deformation $\Delta\mathbf{a}$ is omitted for brevity.
3.2 MoVieS: Unify Appearance, Geometry and Motion
As illustrated in Figure 1, the framework of the proposed MoVieS provides a unified approach to simultaneously model appearance, geometry and motion of dynamic scenes. It consists of a feature backbone with camera and time information for extracting features from input video frames, followed by dedicated heads for depth, splatter and motion estimation.
@@ 3.2.1 Feature Backbone
Given an input posed video $\mathcal{V}=\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i},t_{i}\}_{i=1}^{N}$, we first patchify the input images $\mathbf{I}_{i}$ and use a pretrained image encoder (oquab2024dinov2, ) to extract their features. To effectively incorporate camera information, we adopt two complementary strategies to embed the camera parameters into image features: (1) Plücker embedding: camera pose $\mathbf{P}_{i}$ and intrinsics $\mathbf{K}_{i}$ are encoded into pixel-aligned Plücker embeddings (sitzmann2021light, ), which are then downsampled and fused with image features via spatial-wise addition; (2) Camera token: both $\mathbf{P}_{i}$ and $\mathbf{K}_{i}$ are passed through a linear layer to generate a camera token, which is appended to the sequence of image tokens. Plücker embedding provides a dense and spatially aligned encoding of camera geometry, while the camera token interacts globally with image tokens through attention, enabling more holistic feature reasoning. Ablation study on these two camera injection manners is validated in Section 4.4.
To inform the model that input images originate from a temporally ordered video, we additionally encode each timestamp $t_{i}\in[0,1]$ using sinusoidal positional encoding (mildenhall2020nerf, ) to produce a time token, which is then concatenated with the aforementioned image and camera tokens.
After tokenizing input images, camera parameters, and timestamps, we apply geometrically pretrained attention blocks (vaswani2017attention, ) from VGGT (wang2025vggt, ) to enable interactions among image tokens across video frames. This produces a set of shared feature tokens enriched with inter-frame context as well as camera and temporal information, which are then used for predicting various properties of dynamic scenes.

### v2

@@ Abstract
We present MoVieS, a Motion-aware View Synthesis model that reconstructs 4D dynamic scenes from monocular videos in one second. It represents dynamic 3D scenes with pixel-aligned Gaussian primitives and explicitly supervises their time-varying motions. This allows, for the first time, the unified modeling of appearance, geometry and motion from monocular videos, and enables reconstruction, view synthesis and 3D point tracking within a single learning-based framework. By bridging view synthesis with geometry reconstruction, MoVieS enables large-scale training on diverse datasets with minimal dependence on task-specific supervision. As a result, it also naturally supports a wide range of zero-shot applications, such as scene flow estimation and moving object segmentation. Extensive experiments validate the effectiveness and efficiency of MoVieS across multiple tasks, achieving competitive performance while offering several orders of magnitude speedups.
{NoHyper}: Equal contribution; $\dagger$: Project lead; $\ddagger$: Corresponding author.
@@ 1 Introduction
Humans and animals perceive a continuous stream of observations from a dynamic 3D world, and effortlessly interpret its underlying geometry and motion. Replicating this capability is essential for any embodied agent that must understand and act in the physical world.
@@
Recent advances have made great strides in individual 3D tasks, such as monocular depth estimation [54, 23, 84, 2, 6], 3D scene reconstruction [56, 65, 74, 91, 68], novel view synthesis [45, 12, 24, 76, 43] and point tracking [15, 11, 71, 22, 77]. However, besides treating each task in isolation, most existing view synthesis and reconstruction studies focus on static scenes and require costly per-scene optimization without learning prior knowledge. Real-world environments are inherently dynamic and diverse, and all the aforementioned 3D scene understanding tasks could share common underlying principles.
Figure 1: Overview. MoVieS consists of a shared image encoder, an attention-based feature backbone (Sec. 3.2.1), and three heads (Sec. 3.2.2) to jointly model appearance, geometry and motion. Motion head is time-conditioned to model dynamic content with respect to several query timestamps. Normalized XYZ values in the 3D space of motion maps are treated as RGB channels for visualization. Time-varying Gaussian attributes are omitted and point clouds with color and identity Gaussian attributes are visualized here for brevity.
@@
Motivated by this, we introduce MoVieS, a Motion-aware dynamic View Synthesis model for feed-forward 4D reconstruction of monocular videos, that jointly models scenes’ appearance, geometry and motion. MoVieS represents 3D dynamic scenes using renderable and deformable 3D particles, termed dynamic splatter pixels, and utilizes a differentiable 3D Gaussian rendering framework [24]. Specifically, following recent practices on feed-forward view synthesis [5, 63, 92, 79, 34], each input pixel is mapped to a 3D Gaussian primitive, with its 3D location determined by predicted depth. To model dynamics, MoVieS regresses per-pixel motion displacements toward arbitrary query timestamps, enabling temporal tracking of each splatter pixel. This design facilitates coherent reconstruction of both 3D geometry and appearance across camera viewpoints and temporal frames.
@@
As shown in Figure 1, MoVieS is built upon a large-scale pretrained transformer backbone [47, 68], which encodes each video frame independently and aggregates their information via attentions [66]. The aggregated features are then processed by specialized prediction heads: (1) a depth head estimates depth for each input frame, (2) a splatter head predicts per-pixel 3D Gaussian [24] appearance attributes, such as color and opacity, for novel view rendering, (3) a motion head estimates the time-conditioned movements of Gaussian primitives towards a target timestamp, allowing us to track its temporal evolution.
@@
Benefiting from the unified architecture, MoVieS can be trained on large-scale datasets featuring both static [97, 75, 29] and dynamic [44, 3] scenes, as well as point tracking datasets [96, 21, 20]. During inference, it takes a monocular video, whether depicting a static or dynamic scene, and reconstructs per-pixel 3D Gaussian primitives along with their motion attributes at any target timestamp, enabling view synthesis, depth estimation, and 3D point tracking in a single model.
@@
Extensive experiments on diverse benchmarks [97, 90, 14, 25] reveal that MoVieS achieves competitive performance across a variety of 4D perception tasks, while being several orders of magnitude faster than the existing state of the art. Furthermore, empowered by novel view synthesis as a proxy task, MoVieS enables dense motion understanding from sparse tracking supervision. This naturally gives rise to various applications, such as scene flow estimation and moving object segmentation, further broadening the potential of our approach.
@@
  • We introduce MoVieS, a novel feed-forward framework that jointly models appearance, geometry and motion for 4D scene perception from monocular videos.
@@ 2 Related Work
Figure 2: Motion Head. Given $M$ query timesteps, the proposed motion head is conditioned via adaptive layer normalization (AdaLN) and predicts 3D displacements for each input pixel. After rasterization using the $M$ corresponding query-time cameras, output images in shape $M\times 3\times H\times W$ are rendered for supervision. Gaussian attribute deformation $\Delta\mathbf{a}$ is omitted for brevity.
@@ 2.1 Feed-forward 3D Reconstruction
Traditional 3D reconstruction relies on dense multi-view supervision and per-scene optimization [16, 46, 56, 45]. Recent learning-based methods leverage large-scale priors to enable feed-forward prediction of depth, pose, and other 3D properties [69, 54, 51, 17]. With differentiable rendering techniques such as 3D Gaussian Splatting (3DGS) [24], some works further regress pixel-wise Gaussian attributes for novel view synthesis [5, 9, 92]. DUSt3R [74] pioneered the direct regression of pixel-aligned pointmaps in a canonical space from image pairs. Its extensions [27, 67, 39, 59, 94] generalize this framework with multi-view, streaming, and 3DGS integration. VGGT [68] unifies these advances with a strong image encoder [47] and task-specific heads. However, all of these works are limited to static scenes. In contrast, in this work, we extend feed-forward 3DGS reconstruction to dynamic scenarios with moving objects.
2.2 Dynamic Reconstruction and View Synthesis
Compared to static reconstruction, dynamic reconstruction remains underexplored. Extensions of DUSt3R [8, 61, 20] handle dynamic reconstruction in a plug-and-play manner [8] or leveraging foundation models [91, 42, 87] with monocular depth [2, 50], optical flow [64], and point tracking [22], but remain limited to two-frame inputs and output only sparse point clouds. CUT3R [73] extends to video inputs via recurrent updates, while diffusion-based methods [82, 19] treat reconstruction as conditional generation, though at the cost of multiple passes per sequence.
Instead of geometric structure, dynamic novel view synthesis focuses on rendering quality and view-dependent effects. NeRF-based methods [45, 12, 4] reconstruct dynamic scenes from multi-view [1, 70, 28] or monocular videos [52, 48, 32, 95], while 3DGS [24] extends this with 4D primitives [85, 30] or deformable fields [76, 43, 86, 35, 37]. Recent works further express Gaussian motion explicitly [72, 62, 60, 26, 36], but require scratch training, iterative optimization, and external supervision from point tracking or optical flow estimation [11, 64].
@@ 2.2 Dynamic Reconstruction and View Synthesis
For feed-forward 4D view synthesis, STORM [83] is designed for outdoor driving scenes from multi-view videos. BTimer [33] and NutWorld [57] estimate 3DGS attributes from monocular inputs. However, BTimer predicts independent pixel-wise 3D Gaussians per timestamp without modeling their frame-wise relations and needs an extra enhancer module for smooth intermediate frames. NutWorld models Gaussian motion but lacks explicit supervision, relying heavily on pretrained depth [6] and flow [80] estimation models, and uses an orthographic camera, which further leads to projection distortions.
@@ 3 Method
Following previous studies on monocular 4D reconstruction [95, 72, 37, 26, 33, 83], given a posed video $\mathcal{V}\coloneqq\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i},t_{i}\}_{i=1}^{N}$ with frames $\mathbf{I}_{i}\in\mathbb{R}^{3\times H\times W}$, camera poses $\mathbf{P}_{i}\in\mathbb{SE}(3)$, intrinsics $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ and timestamps $t_{i}\in[0,1]$, we aim to train a generic model that jointly reconstructs its appearance, geometry, and motion. We leave the handling of unposed videos to future work.
@@ 3.1 Dynamic Splatter Pixel
To model 3D scenes with moving contents, we propose a novel representation, namely dynamic splatter pixel, which decomposes dynamic scenes into a set of static Gaussian primitives and their corresponding deformation fields. Given an input video $\mathcal{V}$, each pixel in the $i$-th frame $\mathbf{I}_{i}$ is associated with a splatter pixel $\mathbf{g}_{i}$ [5, 63, 92] in a shared canonical space of the first frame camera’s coordinate system. Each $\mathbf{g}\coloneqq\{\mathbf{x},\mathbf{a}\}$ is parameterized by its position $\mathbf{x}\in\mathbb{R}^{3}$ in the canonical space and other rendering attributes $\mathbf{a}\in\mathbb{R}^{11}$, including rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$, scale $\mathbf{s}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, and color $\mathbf{c}_{\text{rgb}}\in\mathbb{R}^{3}$ [24]. Considering splatter pixels were originally designed for static scenes, we decouple motion from geometric structure to adapt them for dynamic scenes. An additional time-dependent deformation field is introduced, in which each splatter pixel $\mathbf{g}$ is associated with $\mathbf{m}(t)\coloneqq\{\Delta\mathbf{x}(t),\Delta\mathbf{a}(t)\}$. $\Delta\mathbf{x}(t)\in\mathbb{R}^{3}$ is the motion vector of a splatter pixel at time $t$ with respect to the canonical 3D space, and $\Delta\mathbf{a}(t)$ is the change of the corresponding attributes of the splatter pixel at time $t$. Therefore, the splatter pixel $\mathbf{g}$ is deformed to time $t$ as:
@@
By combining static splatter pixels and their deformation fields, we establish the correspondence between each Gaussian primitive and its temporal dynamics, thereby enabling dynamic scene modeling and dense motion estimation.
3.2 Unify Appearance, Geometry and Motion
@@ 3.2.1 Feature Backbone
Given a posed video $\mathcal{V}=\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i},t_{i}\}_{i=1}^{N}$, we first patchify each input image $\mathbf{I}_{i}$ and use a pretrained image encoder [47] to extract their features as shown in Figure 1. To effectively incorporate camera information, we adopt two complementary strategies to embed the camera parameters into image features: (1) Plücker embedding: camera pose $\mathbf{P}_{i}$ and intrinsics $\mathbf{K}_{i}$ are transformed into pixel-aligned Plücker embeddings [58], which are then downsampled and fused with image features via spatial-wise addition; (2) Camera token: $\mathbf{P}_{i}$ and $\mathbf{K}_{i}$ are passed through a linear layer and encoded to a camera token, which is appended to the sequence of image tokens. Ablation study on these two camera injection manners is validated in Sec. 4.4.
To inform the model that input images originate from a temporally ordered video, we additionally encode the timestamp $t_{i}\in[0,1]$ of input frames by sinusoidal positional encoding [45] to produce a timestamp token, which is then concatenated with the aforementioned image and camera tokens as shown in Figure 1. After tokenizing input images, camera parameters, and timestamps, we apply the geometrically pretrained attention blocks [66] from VGGT [68] to enable interactions among image tokens across video frames. This produces a set of shared feature tokens enriched with inter-frame context as well as camera and temporal information, which are then used for predicting various properties of dynamic scenes. The effect of VGGT initialization is discussed in Sec. 4.4.